# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v1

@@ III Model setup and analysis details
We have investigated four physics models that have been used to simulate heavy ion collisions. Three of them are dynamical models (AVFD, AMPT, EPOS) and one is a parameterization model (HYDJET++). The AVFD model is a model designed to study CME-related physics and has the capability to implement axial charge current and magnetic field. The other three models do not have those CME capabilities and are useful only for background studies which are the main purpose of this work. The centralities of the models are determined by the charged hadron multiplicity within $|\eta|<0.5$ as in experiments [35].
For further insights, we have also investigate background behaviors with two versions of a toy model. The inputs to the toy models are parameterizations of measured data.
Figure 2: The $m_{\rm inv}$ distributions. Shown are the OS pair $m_{\rm inv}$ distribution subtracted by the SS pair one from AVFD with axial current densities $n_{5}/s=0,0.1$, and $0.2$ (upper left), AMPT (upper center), EPOS4 (upper right), HYDJET++ (lower left), Toy Model I simulation including primordial pions, and $K_{S}$ and $\rho^{0}$ resonance-decay pions (lower center), and Toy Model II simulation including in addition decay pions from a mass continuum (lower right). Particles used to form the $m_{\rm inv}$ distributions are taken from all charged hadrons (defined to be $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in our study and treated as pions with pion mass) and are from the kinematic range of $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$. In the toy models all final-state particles are charged pions.
@@ III.1 AVFD Model
The AVFD model [36, 37, 38] is an anomalous fluid dynamics developed to describe the evolution of chiral fermion currents in the quark-gluon plasma (QGP) created in relativistic heavy ion collisions in addition to the normal VISHNU simulations [39]. The average magnetic field is calculated from the spectator protons and it is nominally along the direction perpendicular to the RP. In the model, the magnetic field is quantified by event-by-event simulations [40] taking into account the fluctuations of the magnetic field direction with respect to the RP. A modest time evolution is assumed for the decreasing magnetic field with a typical lifetime comparable to the initial time of hydrodynamic evolution $\sim 0.6\text{ fm}/c$. The initial condition for the axial charge density ($n_{5}$) is dynamically generated in AVFD to be proportional to the entropy density ($s$), and the strength is set via the proportionality coefficient $n_{5}/s$. Three $n_{5}/s$ values are used in the simulations, $n_{5}/s$=0, 0.1, and 0.2. The QGP medium evolution is simulated by viscous hydrodynamics to describe the bulk background in heavy ion collisions, with transport parameters for the diffusion coefficient as well as the relaxation time. A final hadronic stage after the hydrodynamic freeze-out is included with hadronic re-scatterings and resonance decays.
Au+Au collisions are simulated by AVFD at $\sqrt{s_{{}_{NN}}}=200$ GeV. In this study we focus on the centrality range of 30–40% [28]. Figure 2 (upper left) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for the three values of $n_{5}/s$. The CME implemented in AVFD affects the relative OS and SS pair $m_{\rm inv}$ distributions–the distribution flattens with larger $n_{5}/s$. This arises from contributions from back-to-back OS pairs (hence larger $m_{\rm inv}$) and near-side SS pairs (hence smaller $m_{\rm inv}$) as a result of the CME, resulting in an enhancement at large $m_{\rm inv}$ and depletion at small $m_{\rm inv}$ for large $n_{5}/s$ compared to small one. Several resonance peaks are apparent in the $m_{\rm inv}$ distributions, $\rho^{0}$, $f_{0}(980)$, and $f_{2}(1270)$. No intrinsic mass width is implemented in AVFD, so the resonances appear as $\delta$-functions. Weak-decay hadrons are treated as stable particles in AVFD, so no $K_{S}$ peak is present in the $m_{\rm inv}$ distribution.
@@ III.3 EPOS4 Model
EPOS4 [45] is a Monte Carlo model to simulate high-energy proton-proton and nucleus-nucleus collisions, among others. At high energies many interactions (nucleonic or partonic) happen simutaneously at once, not sequentially. The parallel multi-parton interactions are modeled via the Pomeron exchange mechanism and result in complex configurations composed of many strings [46, 47]. The string decay products further interact, forming a medium treated by the core-corona picture followed by fluid dynamic evolution of the former and ordinary hadronic interactions in the latter [48]. The EPOS4 model has been extensively tested and tuned by proton-proton and heavy ion data.
MB Au+Au collisions are simulated by EPOS4 (version 4.0.0) at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. Figure 2 (upper right) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $K_{S}$ and $\rho^{0}$ resonance peaks are evident. No intrinsic mass width is implemented in EPOS, so the resonance peaks are sharp $\delta$-functions. All $K_{S}$ particles are decayed in EPOS and their decay pions are treated as “primordial” POIs.
@@ III.4 HYDJET++ Model
HYDJET++ is an event generator to simulate heavy ion collisions by combining two independent components, namely, the soft physics part and the hard physics part [49, 50]. The former is determined by thermal equilibrium where hadrons are produced on the hypersurface represented by a parameterization of relativistic hydrodynamics with given freeze-out conditions. Chemical and kinetic freeze-outs are separate, and hadronic resonances are decayed. The hard physics part starts with an initial parton configuration from PYTHIA [51], lets partons rescatter traversing the hot and dense nuclear medium, and losing energy via collisions and radiative gluon emission, and follows with parton hadronization and particle formation. The model parameters are tuned to reproduce heavy ion data on charged particle multiplicity, $p_{\rm T}$-spectra and flow.
MB Au+Au collisions are generated by HYDJET++ (version 2.4) at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. Figure 2 (lower left) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $K_{S}$, $\rho^{0}$, and $f_{0}(980)$ resonance peaks are evident. No intrinsic mass width is implemented in HYDJET, so the resonance peaks are sharp $\delta$-functions. All $K_{S}$ particles are decayed in HYDJET and their decay pions are treated as “primordial” POIs.
@@ III.5 Toy Model I
To gain insights, we have also generated particles using a toy model [34]. Toy models are useful and convenient to investigate the outcome of an analysis method given the known input to the model.
Two versions of the toy model are examined. In Toy Model I, we include primordial pions and two resonances, $K_{S}$ and $\rho^{0}$. The resonance mass is generated according to the Breit-Wignar distribution of proper mass and width [52]. The primordial pion, $K_{S}$, and $\rho^{0}$ input information are taken from experimental measurements of Au+Au collisions at 200 GeV in the centrality range of 30–40%, which we currently focus on. We generate particles uniformly within rapidity range of $-1.5<y<1.5$, taking event-by-event particle multiplicities according to Poisson statistics about the means of three times the measured average densities [53, 35, 54]. The $p_{\rm T}$ spectra are parameterized based on the measured data [55, 56], namely,
@@
where transverse momentum $m_{\rm T}\equiv\sqrt{p_{\rm T}^{2}+m_{0}^{2}}$ with $m_{0}$ being the rest mass of the corresponding particle. The particle or resonance $p_{\rm T}$ is sampled from Eq. 11. At a given $p_{\rm T}$, the average $\left\langle v_{2}\right\rangle$ of pion, $K_{S}$, or $\rho^{0}$ meson is calculate based the following parameterization of data motivated by the number-of-constituent-quark scaling [57],
$\left\langle v_{2}\right\rangle(p_{\rm T})=\frac{2a}{1+\exp\left(-\frac{(m_{\rm T}-m_{0})/2-b}{c}\right)}-2d\,.$ (12)
A 40% $v_{2}$ fluctuations are included about the parameterized $\left\langle v_{2}\right\rangle(p_{\rm T})$ event-by-event to obtain the final $v_{2}$ of the particle, based on which the azimuthal angle is generated about the fixed event plane at $\psi=0$. All parameters used in the toy model can be found in Ref. [34].
Toy model data are generated, using parameters corresponding to the 30–40% centrality of Au+Au collisions at 200 GeV. Figure 2 (lower center) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs. We require all resonances in our toy model (i.e. $K_{S}$ and $\rho^{0}$) to decay into a pair of charged pions. The decay is isotropic in the parent rest frame, and the decay pions are properly boosted to the lab frame. The $K_{S}$ and $\rho^{0}$ resonance peaks are evident.
@@ III.6 Toy Model II
As shown in Fig. 2, all physics models have a mass continuum in the $m_{\rm inv}$ distribution on which the resonance peaks reside. To mimic this, we add a mass continuum in our second version of the toy model. The mass distribution of the continuum is generated according to a probability $\propto e^{-m_{\rm inv}c^{2}/{\rm GeV}}$ between the two-pion mass threshold (0.28 GeV/$c^{2}$) and 5 GeV/$c^{2}$. Given a generated mass value, a Breit-Wignar distribution (mean at the mass value and width equal to the $\rho^{0}$ width [52]) is sampled to obtain the mass of the continuum “resonance”. All particles in the generated mass continuum are decayed into $\pi^{+}\pi^{-}$ isotropically in the parent rest frame. The decay pions are boosted properly to the lab frame.
@@ III.7 Analysis Details
POIs are taken to be all charged hadrons, defined to be $\pi^{\pm}$, $K^{\pm}$, $p$ and $\bar{p}$ in our study, and are within the transverse momentum range of $20.2<p_{\rm T}<$ 2 GeV/$c$ and a certain central pseudorapidity ($\eta$) range, typical of midrapidity detectors such as the STAR experiment [58].
In the ESS method, the POIs are taken to be within $|\eta|<1$ (or $|\eta|<2$). The event selection quantity pair $\hat{q}_{2,{\rm pair}}^{2}$ is computed by the two-particle cumulant method in Eqs. 5,9, always using all particle pairs of the same POIs.
@@
Figure 4 shows the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ and $\hat{q}_{2}^{2}$ distributions from the ESS and ESE methods, respectively. The two distributions are similar for each of the models; the $\hat{q}_{2}^{2}$ distribution ($|\eta|<0.3$) is broader than the $\hat{q}_{2,{\rm pair}}^{2}$ distribution ($|\eta|<1$) because of larger fluctuations in the smaller acceptance. We have checked that the distribution of the single particle $\hat{q}_{2}^{2}$ calculated from $|\eta|<1$ is similar to the pair $\hat{q}_{2,{\rm pair}}^{2}$ from the same acceptance.
Figure 3: The $\hat{q}_{2}^{2}$ distributions. Shown are the single $\hat{q}_{2}^{2}$ and pair $\hat{q}_{2,{\rm pair}}^{2}$ distributions of POIs from AVFD with $n_{5}/s=0$ (upper left), AMPT (upper center), EPOS (upper right), HYDJET (lower left), Toy Model I (lower center), and Toy Model II (lower right). Shown in all panels are for the 30–40% centrality of Au+Au collisions, as examples. The $\hat{q}_{2}^{2}$ (used in the ESE analysis) is calculated from particles in $|\eta|<0.3$, and the pair $\hat{q}_{2,{\rm pair}}^{2}$ (used in the ESS analysis) is calculated from particles within $|\eta|<1$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$.
@@
Figure 4: The average $v_{2}^{2}\{2\}$. Shown are the two-particle cumulant $v_{2}^{2}\{2\}$ of POIs and $v_{2,{\rm pair}}^{2}\{2\}$ of particle pairs of POIs as functions of centrality from AVFD for $n_{5}/s=0$ (upper left), AMPT (upper center), EPOS (upper right), HYDJET (lower left), Toy Model I (lower center), and Toy Model II (lower right). POIs are from acceptances $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$. The large departure of $v_{2,{\rm pair}}^{2}\{2\}$ from $v_{2}^{2}\{2\}$, both calculated by the cumulant method (Eq. 8), is due to self-correlations where two pairs share a common particle. The issue is severe in peripheral collisions except for HYDJET where the average multiplicity is significantly larger than the rest models; for reference, for 30–40% centrality of Au+Au, the average multiplicities are 151, 169, 88, and 255 for AVFD, AMPT, EPOS, and HYDJET, respectively.
@@
The $\hat{q}_{2}^{2}$ variable calculated from single particles has been widely used. The $\hat{q}_{2,{\rm pair}}^{2}$ variable calculated from particle pairs is relatively new. To gain insights, we calculate by Eq. 8 the two-particle cumulant $v_{2}\{2\}$ of single particles of the POIs in events grouped by $\hat{q}_{2,{\rm pair}}^{2}$. Such a quantity can also be calculated for particle pairs of POIs, in which case the $\phi_{1}$ and $\phi_{2}$ in Eq. 8 are the azimuthal angles of two pairs, each formed by two POIs. Figure 4 shows $v_{2}\{2\}$ (from 2 POIs) and $v_{2,{\rm pair}}\{2\}$ (from 2 pairs of POIs) as functions of centrality in different simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200\text{ GeV}$. They roughly agree at high multiplicity (HYDJET, or central collisions of other models), and $v_{2,{\rm pair}}\{2\}$ becomes much larger than $v_{2}\{2\}$ at low multiplicity (peripheral collisions of models except for HYDJET). This discrepancy comes from the autocorrelations where the two pairs share one common POI. To get a feeling of this effect, consider a simplified scenario with $\phi_{{\rm pair}}=(\phi_{1}+\phi_{2})/2$, which is a statistically unbiased assumption. For two pairs with indices $(1,3)$ and $(2,3)$ where particle 3 is shared, their correlation is $\cos 2[(\phi_{1}+\phi_{3})/2-(\phi_{2}+\phi_{3})/2]=\cos(\phi_{1}-\phi_{2})$, and the fraction of this case in the $v_{2,{\rm pair}}^{2}\{2\}$ calculation is the triplet multiplicity over the quadruplet one, which is $\sim 1/N$. Therefore, besides the common flow component, $v_{2,{\rm pair}}\{2\}$ has an extra part $\langle\cos(\phi_{1}-\phi_{2})\rangle/N$, which accounts qualitatively for its excess over $v_{2}\{2\}$ when multiplicity is low.
@@
Likewise, in each event class according to $\hat{q}_{2}^{2}$ in the ESE analysis, we also use the known EP of $\psi=0$ in models to calculate $\left\langle v_{2}\right\rangle$ by Eq. 10 and $\Delta\gamma$ by Eqs. 2,3. Here, the POI is taken from the west or east subevent in calculating $\left\langle v_{2}\right\rangle$, and in calculating $\Delta\gamma$, both POIs are taken from either the west subevent or the east subevent. In addition, we use the two-particle cumulant (Eq. 8) to compute $v_{2}\{2\}$ where one particle is taken from the west subevent and the other from the east subevent, and the three-particle correlator [7] to compute

### v3

@@ III Model setup and analysis details
We investigate four physics models that have been used to simulate heavy ion collisions. Three of them are dynamical models (avfd, ampt, epos4) and the fourth is a parameterization model (hydjet++). The avfd model is designed to study CME-related physics and has the capability to implement axial charge current and magnetic field. The other three models do not have those CME capabilities and are useful only for background studies, which is the main purpose of this work. The centralities of the models are determined by the charged hadron multiplicity within $|\eta|<0.5$ as in the STAR experiments [45].
For further insights, we also investigate background behaviors with two versions of a toy model. The inputs to the toy models are parameterizations of the measured data.
Figure 2: $m_{\rm inv}$ distributions. Shown are the per-event $m_{\rm inv}$ distribution of OS pairs subtracted by that of SS pairs from avfd with axial current densities $n_{5}/s=0,0.1$, and $0.2$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I simulation including primordial pions, and $K_{S}$ and $\rho^{0}$-resonance decay pions (lower center), and Toy Model II simulation including in addition decay pions from a mass continuum (lower right). Particles used to form the $m_{\rm inv}$ distributions are taken from all charged hadrons (defined to be $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in our study and treated as pions with pion mass) and are from the kinematic range of $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$. In the toy models all final-state particles are charged pions.
@@ III.1 AVFD Model
The avfd model [46, 47, 48] is an anomalous fluid dynamics developed to describe the evolution of chiral fermion currents in the quark-gluon plasma (QGP) created in relativistic heavy ion collisions in addition to the normal VISHNU simulations [49]. The average magnetic field is calculated from the spectator protons, and it is nominally along the direction perpendicular to the RP. In the model, the magnetic field is quantified by event-by-event simulations [50] taking into account the fluctuations of the magnetic field direction with respect to the RP. A modest time evolution is assumed for the decreasing magnetic field with a typical lifetime comparable to the initial time of hydrodynamic evolution $\sim 0.6\text{ fm}/c$. The initial condition for the axial charge density ($n_{5}$) is dynamically generated in avfd to be proportional to the entropy density ($s$), and the strength is set via the proportionality coefficient $n_{5}/s$. Three $n_{5}/s$ values are used in the simulations, $n_{5}/s$=0, 0.1, and 0.2. The QGP medium evolution is simulated by viscous hydrodynamics to describe the bulk background in heavy ion collisions, with transport parameters for the diffusion coefficient as well as the relaxation time. A final hadronic stage after the hydrodynamic freeze-out is included with hadronic re-scatterings and resonance decays.
Au+Au collisions are simulated by avfd at $\sqrt{s_{{}_{NN}}}=200$ GeV in [34]. In this study, we focus on the centrality range of 30–40% as in [34]. Figure 2 (upper left) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for the three values of $n_{5}/s$. The CME implemented in avfd affects the relative OS and SS pair $m_{\rm inv}$ distributions–the distribution flattens with larger $n_{5}/s$. This arises from contributions from back-to-back OS pairs (hence larger $m_{\rm inv}$) and near-side SS pairs (hence smaller $m_{\rm inv}$) as a result of the CME, resulting in enhancement at large $m_{\rm inv}$ and depletion at small $m_{\rm inv}$ for large $n_{5}/s$ compared to small one. Several resonance peaks are apparent in the $m_{\rm inv}$ distributions: $\rho^{0}$, $f_{0}(980)$, and $f_{2}(1270)$. No intrinsic mass width is implemented in avfd, so the resonances appear as $\delta$-functions. Weak-decay hadrons are treated as stable particles in avfd, so no $K_{S}$ peak is present in the $m_{\rm inv}$ distribution.
@@ III.3 EPOS4 Model
epos4 [55] is a Monte Carlo model to simulate high-energy proton-proton and nucleus-nucleus collisions, among others. At high energies many (nucleonic and partonic) interactions happen simultaneously at once, not sequentially. Parallel multiparton interactions are modeled via the pomeron exchange mechanism and result in complex configurations composed of many strings [56, 57]. The string decay products further interact, forming a medium treated by the core-corona picture followed by fluid dynamic evolution of the former and ordinary hadronic interactions in the latter [58]. The epos4 model has been extensively tested and tuned using proton-proton and heavy ion data.
MB Au+Au collisions are simulated by epos4 (version 4.0.0) at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. Figure 2 (upper right) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $K_{S}$ and $\rho^{0}$ resonance peaks are evident. No intrinsic mass width is implemented in epos4, so the resonance peaks are sharp $\delta$-functions. All $K_{S}$ particles are decayed in epos4 and their decay pions are treated as “primordial” pions.
@@ III.4 HYDJET++ Model
hydjet++ is an event generator to simulate heavy ion collisions by combining two independent components, namely, the soft physics part and the hard physics part [59, 60]. The former is determined by thermal equilibrium where hadrons are produced on the hypersurface represented by a parameterization of relativistic hydrodynamics with given freeze-out conditions. Chemical and kinetic freeze-outs are separate, and hadronic resonances are decayed. The hard physics part starts with an initial parton configuration from pythia [61], lets partons rescatter traversing the hot and dense nuclear medium and losing energy via collisions and radiative gluon emission, and follows with parton hadronization and particle formation. The model parameters are tuned to reproduce heavy ion data on charged particle multiplicity, $p_{\rm T}$-spectra and flow.
MB Au+Au collisions are generated by hydjet++ (version 2.4) at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. Figure 2 (lower left) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $K_{S}$, $\rho^{0}$, and $f_{0}(980)$ resonance peaks are evident. No intrinsic mass width is implemented in hydjet++  so the resonance peaks are sharp $\delta$-functions. All $K_{S}$ particles are decayed in hydjet++ and their decay pions are treated as “primordial” pions.
@@ III.5 Toy Model I
To gain insights, we have also generated particles using a toy model [44]. Toy models are useful and convenient for investigating the outcome of an analysis method given the known input to the model.
Two versions of the toy model are examined. In Toy Model I, we include primordial pions and two resonances, $K_{S}$ and $\rho^{0}$. The resonance mass is generated according to the Breit-Wignar distribution of proper mass and width [62]. We generate particles uniformly within the rapidity range of $-1.5<y<1.5$, taking event-by-event particle multiplicities according to Poisson statistics about the means of three times the measured average pseudorapidity densities [63, 45, 64]. The $p_{\rm T}$ spectra are parameterized based on the measured data [65, 66], namely,
@@
where transverse momentum $m_{\rm T}\equiv\sqrt{p_{\rm T}^{2}+m_{0}^{2}}$ with $m_{0}$ being the rest mass of the corresponding particle. The particle or resonance $p_{\rm T}$ is sampled from Eq. 11. At a given $p_{\rm T}$, the average $\left\langle v_{2}\right\rangle$ of pion, $K_{S}$, or $\rho^{0}$ meson is calculated based on the following parameterization of data motivated by the number-of-constituent-quark scaling [67],
$\left\langle v_{2}\right\rangle(p_{\rm T})=n_{q}\left[\frac{a}{1+\exp\left(-\frac{(m_{\rm T}-m_{0})/n_{q}-b}{c}\right)}-d\right]\,,$ (12)
where $n_{q}$ is the number of constituent quarks for a given particle and, for our purpose, $n_{q}=2$. A 40% $v_{2}$ fluctuation is included about the parameterized $\left\langle v_{2}\right\rangle(p_{\rm T})$ event-by-event to obtain the final $v_{2}$ of the particle, based on which the azimuthal angle is generated about the fixed event plane at $\psi=0$.
The primordial pion, $K_{S}$, and $\rho^{0}$ input information are taken from experimental measurements of Au+Au collisions at 200 GeV in the centrality range of 30–40%, which we currently focus on. All parameters used in the toy model can be found in Ref. [44]. Figure 2 (lower center) shows the difference of the $m_{\rm inv}$ distributions between the OS and SS pairs from the generated toy-model data. We require that all resonances in our toy model (i.e. $K_{S}$ and $\rho^{0}$) to decay into a pair of charged pions. The decay is isotropic in the parent rest frame, and the decay pions are properly boosted to the lab frame. The $K_{S}$ and $\rho^{0}$ resonance peaks are evident.
@@ III.6 Toy Model II
As shown in Fig. 2, all physics models have a mass continuum in the $m_{\rm inv}$ distribution on which the resonance peaks reside. To mimic this, we add a mass continuum in our second version of the toy model. The mass distribution of the continuum is generated according to a probability $\propto e^{-m_{\rm inv}c^{2}/{\rm GeV}}$ between the two-pion mass threshold (0.28 GeV/$c^{2}$) and 5 GeV/$c^{2}$. Given a generated mass value, a Breit-Wignar distribution (mean at the mass value and width equal to the $\rho^{0}$ width [62]) is sampled to obtain the mass of the continuum “resonance.” All particles in the generated mass continuum are decayed into $\pi^{+}\pi^{-}$ isotropically in the parent rest frame. The decay pions are properly boosted to the lab frame.
@@ III.7 Analysis Details
Figure 3: $\hat{q}_{2}^{2}$ distributions. Shown are the single $\hat{q}_{2}^{2}$ and pair $\hat{q}_{2,{\rm pair}}^{2}$ distributions of POIs from avfd with $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). Shown in all panels are for the 30–40% centrality of Au+Au collisions, as examples. The $q_{2}^{2}$ (used in the ESE analysis) is calculated from particles in $|\eta|<0.3$, and the pair $q_{2,{\rm pair}}^{2}$ (used in the ESS analysis) is calculated from particles within $|\eta|<1$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$.
POIs are assumed to be all charged hadrons, defined to be $\pi^{\pm}$, $K^{\pm}$, $p$ and $\bar{p}$ in our study, and are within the transverse momentum range of $20.2<p_{\rm T}<$ 2 GeV/$c$ and a certain central pseudorapidity ($\eta$) range, typical of midrapidity detectors such as the STAR experiment [68].
In the ESS method, the POIs are taken to be within $|\eta|<1$ (or $|\eta|<2$). The event selection quantity, pair $\hat{q}_{2,{\rm pair}}^{2}$, is computed by the two-particle cumulant method in Eqs. 5,9, always using all particle pairs of the same sample of POIs.
@@
Figure 3 shows the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ and $\hat{q}_{2}^{2}$ distributions from the ESS and ESE methods, respectively. The two distributions are similar for each of the models; the $\hat{q}_{2}^{2}$ distribution ($|\eta|<0.3$) is broader than the $\hat{q}_{2,{\rm pair}}^{2}$ distribution ($|\eta|<1$), presumably due to larger fluctuations in the smaller acceptance. We have checked that the distribution of the single particle $\hat{q}_{2}^{2}$ calculated from $|\eta|<1$ is similar to the pair $\hat{q}_{2,{\rm pair}}^{2}$ from the same acceptance.
@@
Figure 4: Average $v_{2}^{2}\{2\}$. Shown are the two-particle cumulant $v_{2}^{2}\{2\}$ of POIs and $v_{2,{\rm pair}}^{2}\{2\}$ of particle pairs of POIs as functions of centrality from avfd for $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). POIs are from acceptances $|\eta|<1$ (except for epos4, $|\eta|<2$) and $20.2<p_{\rm T}<$ 2 GeV/$c$. The large departure of $v_{2,{\rm pair}}^{2}\{2\}$ from $v_{2}^{2}\{2\}$, both calculated by the cumulant method (Eq. 8), is largely due to self-correlations where two pairs share a common particle. The issue is severe in peripheral collisions except for hydjet++ where the average multiplicity is significantly larger than the rest models; for reference, for 30–40% centrality of Au+Au, the average multiplicities within $|\eta|<0.5$ are 151, 169, 88, and 255 for avfd, ampt, epos4, and hydjet++, respectively.
@@
The $\hat{q}_{2}^{2}$ variable calculated from single particles has been widely used. The $\hat{q}_{2,{\rm pair}}^{2}$ variable calculated from particle pairs is relatively new [36]. As seen from Eq. 7, the average $\left\langle q_{2}^{2}\right\rangle$ is related to the two-particle cumulant $v_{2}\{2\}$. To gain insights, we calculate by Eq. 8 the $v_{2}\{2\}$ of single POIs and the $v_{2,{\rm pair}}\{2\}$ of particle pairs of POIs; In the latter case, the $\phi_{1}$ and $\phi_{2}$ in Eq. 8 are the azimuthal angles of two pairs, each formed by two POIs. The $v_{2,{\rm pair}}\{2\}$ consists of a component from pairing of “random” (hydrodynamic) particle pairs or a hydrodynamic pair with a resonance-decay pair, and a component from two-resonance $v_{2}\{2\}$. The first component reflects, in a non-trivial way, the single particle $v_{2}$ [69]. The last component is simply resonance (background contributing source) $v_{2}\{2\}$. Both $v_{2}\{2\}$ and $v_{2,{\rm pair}}\{2\}$ contain nonflow contamination, for example, from resonance decays (in the case of $v_{2,{\rm pair}}\{2\}$, the two pairs both contain daughters from the same resonance decay). Figure 4 shows $v_{2}\{2\}$ (from 2 POIs) and $v_{2,{\rm pair}}\{2\}$ (from 2 pairs of POIs) as functions of centrality in different simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200\text{ GeV}$. They roughly agree at high multiplicity (hydjet++, or central collisions of other models), and $v_{2,{\rm pair}}\{2\}$ becomes much larger than $v_{2}\{2\}$ at low multiplicity (peripheral collisions of models except for hydjet++). This discrepancy comes from the self-correlations where the two pairs share one common POI. To get a feeling of this effect, consider a simplified scenario with $\phi_{{\rm pair}}=(\phi_{1}+\phi_{2})/2$. For two pairs with indices $(1,3)$ and $(2,3)$ where particle 3 is shared, their correlation is $\cos 2[(\phi_{1}+\phi_{3})/2-(\phi_{2}+\phi_{3})/2]=\cos(\phi_{1}-\phi_{2})$, and the fraction of this case in the $v_{2,{\rm pair}}^{2}\{2\}$ calculation is the triplet multiplicity over the quadruplet one, which is $\sim 1/N$. Therefore, besides the common flow component, $v_{2,{\rm pair}}\{2\}$ has an extra part $\sim\langle\cos(\phi_{1}-\phi_{2})\rangle/N$, which accounts qualitatively for its excess over $v_{2}\{2\}$ when multiplicity is low.
@@
Likewise, in each event class according to $q_{2}$ in the ESE analysis, we also use the known EP of $\psi=0$ in models to calculate $\left\langle v_{2}\right\rangle$ by Eq. 10 and $\Delta\gamma$ by Eqs. 2,3. Here, the POI is taken from the west or east subevent in calculating $\left\langle v_{2}\right\rangle$, and in calculating $\Delta\gamma$, both POIs are taken from either the west subevent or the east subevent. In addition, we use the two-particle cumulant (Eq. 8) to compute $v_{2}\{2\}$ where one particle is taken from the west subevent and the other from the east subevent, and the three-particle correlator [8] to compute